OpenStudy (anonymous):

Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS

OpenStudy (anonymous):

any ideas yet

OpenStudy (dape):

Start with a contradiction, for example 0=1. Now you can do a lot, for example add x to both sides: x=x+1 and another step, for example divide by 2: x/2 = (x+1)/2 This is our equation, if we do the steps backward (multiply by 2, subtract x), we get to our contradiction.

OpenStudy (anonymous):

u sure

OpenStudy (dape):

Yeah.

OpenStudy (dape):

If you don't believe me, just try to find a x that satisfies x/2 = (x+1)/2 ;)

OpenStudy (anonymous):

lol i believe you. i dont feel like doing work by myself. :)

OpenStudy (dape):

You could also argue that they are the two lines y=x/2 y=(x+1)/2=x/2+1/2, which are parallell. Since two lines never meet there are no solutions.

OpenStudy (anonymous):

would that be anothre example or the same one

OpenStudy (anonymous):

and thanks for helping me :)

OpenStudy (dape):

That's the same example. If you really wanna show off you could do something like \[ \frac{1}{x+5} = 0 \] and say that's it's asymptotic to the x-axis. Then you could argue that it has solutions +∞ and -∞ if you extend the reals to include surreal numbers and be the coolest kid in class.

OpenStudy (anonymous):

lol 1) i dont know what asymptotic means and 2) i dont think that will make me be the coolest kid in my class

OpenStudy (dape):

You must have very lame classmates, everybody knows math is awesome =)

OpenStudy (anonymous):

yeah but i dont take math at school i take online, but my classmates in my other classes might be lame. idk i feel like they are